OpenStudy (anonymous):

Which is an example of the distributive property? A. 9 • (t + 5) = 9t + 45 B. 9 • (t + 5) = (t + 5) • 9 C. 9 • (t + 5) = (9 • t) + 5 D. 9 • (t + 5) = 9t + 5

OpenStudy (anonymous):

Please help!!!!

Vocaloid (vocaloid):

hint: A(B+C) = A*B + A*C 9(t+5) = _____ fill in the blank

OpenStudy (anonymous):

😐

Vocaloid (vocaloid):

just follow the pattern 9(t+5) = 9*t + 9*5 = ?

OpenStudy (anonymous):

My math skills are still at a 3rd grade level man, I only just got fractions.

Vocaloid (vocaloid):

multiply the 9 and the t together multiply the 9 and the 5 together

OpenStudy (anonymous):

Thanks :D

Vocaloid (vocaloid):

well, what did you get as your answer?

OpenStudy (anonymous):

45

OpenStudy (anonymous):

I gived you a medal.

OpenStudy (anonymous):

Chur welkum :)